Bart Incorporated manufactures rotary air fans and uses a production−planning horizon of on minus−half. It allows for regular​ time, part​ time, overtime, and subcontracting production to meet the demand but no more than 12 units may be made by each of these methods. Both anticipation inventory and backorders are allowed. The beginning​ (or current) inventory is 20 units. Their first attempt at a sales and operations plan has resulted in the following.

1

2

3

4

5

6

Total

Inputs

Forecasted demand

100

125

120

80

75

70

570

Workforce Level

80

80

80

80

80

80

480

Part Time

12

12

12

12

12

12

72

Overtime

0

0

0

0

0

0

0

Subcontract

0

0

0

0

0

0

0

Derived

Beginning inventory

20

12

0

0

0

0

Ending inventory

12

0

0

0

0

2

Backlog

0

21

49

37

20

0

127

Calculated

Workforce level​ $5

400

400

400

400

400

400

​$2,400

Part Time​ $6

72

72

72

72

72

72

​$432

Overtime​ $7.5

0

0

0

0

0

0

​$0

Subcontract​ $8

0

0

0

0

0

0

​$0

Inventory holding​ $5

80

30

0

0

0

5

​$115

Back orders​ $20

0

420

980

740

400

0

​$2,540

Total

​$552

​$922

​$1,452

​$1,212

​$872

​$477

​$5,487

What would be the total plan cost if no more than five units each could be made each month by part−​time, ​overtime, and subcontracted​ workers?

A. ​$4,775

B. ​$4,550

C. ​$4,822

D. ​$4,920
Answer #1

If no more than five units each could be made each month by part−​time, ​overtime, and subcontracted​ workers, then the plan will change as below.

1

2

3

4

5

6

Total

Inputs

Forecasted demand

100

125

120

80

75

70

570

Workforce Level

80

80

80

80

80

80

480

Part Time

5

5

5

5 5 5 30

Overtime

Subcontract

Derived

Beginning inventory

20

5

0

0

0

0

Ending inventory

5

0

0

0

0

0

Backlog

0

125-80-5-5=35

120-80-5+35=70

80-80-5+70=65

75-80-5+65=55

70-80-5+55=40

265

Calculated

Workforce level​ $5

400

400

400

400

400

400

​$2,400

Part Time​ $6

6*5=30 30 30 30 30 30 180

Overtime​ $7.5

Subcontract​ $8

Inventory holding​ $5

5*5=25

30

0

0

0

5

​$115

Back orders​ $20

0

20*35=700

20*70=1400 20*65=1300

20*55=1100

20*40=800

​$5300

Total

$​7995


Please give thumbs up if you like my answer. Thanks!